Question 15 TC = 218 +56Q + 302 What is the average variable cost when 12 units are produced? Enter as a value.
Answer #1

Ans

AVC=92

Total cost is give by

TC=218+56Q+3Q^2

we know Total cost consists of Fixed cost and variable cost, if we deduct fixed cost then, we will get variable cost

TC=FC+VC

FC=218

VC=56Q+3Q^2

As, we have to calculate Average variable cost.

Average variable cost(AVC)= VC/Q

=56Q+3Q^2/Q

= Q(56+3Q)/Q

Q and Q get cancelled out, So, AVC=56+3Q

At Q=12

AVC= 56+3*12

AVC= 92